OpenStudy (anonymous):

2v^5 + 12v^3 + 18v dv

OpenStudy (anonymous):

\[\int (2v^5+12v^3 +18v )dv\]?

myininaya (myininaya):

\[\int\limits_{}^{}(2v^5+12v^3+18v )dv=2\frac{v^{5+1}}{5+1}+12\frac{v^{3+1}}{3+1}+18\frac{v^{1+1}}{1+1}\]

OpenStudy (anonymous):

\[1+1\]?

OpenStudy (anonymous):

and a constant:P

myininaya (myininaya):

+C

OpenStudy (anonymous):

the constant is usually not P

OpenStudy (anonymous):

:P

myininaya (myininaya):

nice call elecengineer

OpenStudy (anonymous):

it is usually ;)

OpenStudy (anonymous):

did you get any rest i hope?

myininaya (myininaya):

alittle

OpenStudy (anonymous):

well it is about time for me to turn into a pumpkin myself

myininaya (myininaya):

ok later

OpenStudy (anonymous):

thanks